### v1

# FLARE: Fusing Language Models and Collaborative Architectures for Recommender Enhancement

Abstract
Hybrid recommender systems, combining item IDs and textual descriptions, offer potential for improved accuracy. However, previous work has largely focused on smaller datasets and model architectures. This paper introduces Flare (Fusing Language models and collaborative Architectures for Recommender Enhancement), a novel hybrid recommender that integrates a language model (mT5) with a collaborative filtering model (Bert4Rec) using a Perceiver network. This architecture allows Flare to effectively combine collaborative and content information for enhanced recommendations.
We conduct a two-stage evaluation, first assessing Flare’s performance against established baselines on smaller datasets, where it demonstrates competitive accuracy. Subsequently, we evaluate Flare on a larger, more realistic dataset with a significantly larger item vocabulary, introducing new baselines for this setting. Finally, we showcase Flare’s inherent ability to support critiquing, enabling users to provide feedback and refine recommendations. We further leverage critiquing as an evaluation method to assess the model’s language understanding and its transferability to the recommendation task.
1 Introduction
Sequential recommender systems incorporate the sequence of user interactions to recommend items grounded in the user’s historical preferences and temporal context (Donkers2017-xy; Hidasi2016-qn; sasrec; bert4rec). While early systems modeled sequences of items only, recent work have proposed hybrid recommenders that also model sequential context (e.g. textual descriptions, categories, attributes) for those items (Hidasi2016-zt; Huang2019-pu; Zhang2019-tr). These models suggest that the sequential item context is itself informative, with potential to improve general recommendation quality and handle cold start for new items.
With the recent trend in transformer-based and natural language models, more recent work has focused on recommender models that incorporate free-form text and attributes (recformer; s3rec; carca; u-bert; userllm; tiger). Adding text tends to improve performance over item ID baselines, e.g. Bert4Rec (bert4rec). However, to the best of our knowledge, the effectiveness of hybrid recommenders have been evaluated only on small datasets, specifically ones with a limited number of items. For instance, many models use subsets of the Amazon Product Reviews dataset that contains relatively small item inventory. Examples of common sub-sets include the Video Games (17k items), Office Products (27k) and Instruments (10k).
To investigate the performance of hybrid models on larger datasets, this paper uses two recommenders. The first is a reimplementation of Bert4Rec to serve as an item-only baseline, and Flare, a simple test-bench hybrid recommender that extends Bert4Rec by incorporating a modern text encoder using a Perceiver (perceiver) network. This integration enables the synthesis of text based item context and item embeddings, potentially providing a richer representation for recommendations. Throughout this paper, ‘item-text’ denotes the concatenated key-value representation of all item context expressed in natural language.
When thoroughly tuned, Bert4Rec and Flare models outperform more complex hybrid models on certain datasets. Notably, we observe that on certain datasets, a tuned baseline Bert4Rec model, which uses only IDs, surpasses previously reported hybrid model results that use text. On datasets where Bert4Rec underperforms, our tuned model consistently outperforms the previously reported Bert4Rec baselines used to evaluate hybrid recommenders. These findings suggest that the contribution of adding text may yield less increased performance for next recommendation prediction than previously thought. That said, Flare, which incorporates text, outperforms our tuned Bert4Rec baseline, suggesting that item text improves metrics. In addition to the smaller datasets used by the current benchmarks, we also present new baseline results for Clothing, Shoes and Jewelry dataset from Amazon Product Reviews (376k) which involve larger vocabularies, where we observe similar trends.
Finally, while it is natural to use standard ranking metrics to evaluate hybrid recommenders, we propose critiquing (Chen2012-dv; Gao2021-xo; Jannach2021-bp) as an additional methodology. For hybrid recommenders trained with textual context, we believe that evaluating recommendations with text critiques is a natural way to measure how well the model assimilates item and textual information. To evaluate models with critiquing, we simulate the user providing input for the desired recommendations based on recent recommended items. We present an evaluation methodology for applying critiquing using the Amazon Reviews dataset and establish baselines.
Our contributions include:
  • We propose a conceptually simple hybrid recommender model using an LLM to encode text and a Perceiver network to produce contextual embeddings (Figure 1). These are then combined with ID embeddings for a richer representation.
  • We show that, despite its simplicity, Flare obtains competitive performance on small datasets, approaching state-of-the-art techniques in some cases and outperforming previously reported Bert4Rec baselines used to evaluate hybrid recommenders.
  • Additionally to the smaller datasets used in current benchmarks, we establish new baselines for hybrid models on large vocabulary datasets.
  • We investigate critiquing as a promising methodology to assess and debug performance gains obtained as a result of adding text meta-data. We present a new critiquing-based metric and associated baselines.
2 Related Work
Work in sequence modeling for recommendations have traditionally focused on RNNs (Hidasi2016-qn; Donkers2017-xy), and more recently, using transformers. For example, SASRec (sasrec) proposes using a transformer-based approach to model item sequences for recommendations, using causal self-attention. Bert4Rec (bert4rec) extends SASRec with BERT-style (bert) masked language-model training for item sequences. These seminal models serve as popular baselines in subsequent work.
To improve general recommendation quality, follow-on efforts investigate hybrid models, which augment the sequence of item IDs with a corresponding sequence of item context (e.g. textual descriptions, categories, attributes). For example, HybridBERT4Rec (hybridbert4rec) extends BERT4Rec with a separate tower that models item context, using the set of users that consume that item as input. S3Rec (s3rec) proposes self-supervised pre-training tasks to align items and attributes by utilizing mutual information maximization. CaFe (li2022coarsetofine) explicitly models coarse-grained, or intent-based, sequences along with dense fine-grained sequences to better capture short-term vs long-term preferences. CARCA (carca) combines item sequences, associated context, and item attributes connected via cross-attention to better correlate between older and more recent items in the user profile. U-BERT (u-bert) trains user and content encoders during self-supervised pre-training with content-rich signals, followed by fine-tuning to focus on content-deficient domains. Some efforts explore replacing the explicit item ID representation and rely solely on a contextual representation. For example, Recformer (recformer) proposes an item representation comprising learned embeddings of text-based key-value pairs. TIGER (tiger) investigates a semantic ID representation for items, along with a T5-style (T5) training and generative approach for retrieval.
As described in the next section, Flare shares similarities with some of the models above. It learns embeddings to represent each item, uses a pre-trained text encoder to encode item metadata, and models item sequence with a stack of transformer layers. Flare differs in that it does not require additional pre-training (aside from the “off-the-shelf” text encoder) nor fine-tuning.
Recent efforts explore incorporating collaborative filtering signals directly into a large-language model (LLM), motivated by the intuition that the language semantics encoded in the LLM promote item context interpretation. P5 (p5rec) represents all user–item signals, along with associated context, as natural language sequences that is input to an encoder/decoder model. LC-Rec (lc-rec) uses vector quantization for ID representation and LLM pre-training tasks to align items and language. CLLM4Rec (zhu2023collaborative) adds new user and item tokens to the LLM, along with a training head to model the collaborative filtering signal. Hua et al. (Hua_2023) studies methods to effectively add item IDs to the LLM token vocabulary. User-LLM (userllm) encodes user history into user embeddings, and incorporates them into an LLM via cross-attention. Harte et al. (Harte2023-gq) experiment with extracting item embeddings from a pretrained LLM to make recommendations directly, as well as to initialize Bert4Rec’s item embeddings. Unlike the above, Flare is focused on using encoders and does not explore the use of a language decoder to make recommendations.
We propose critiquing (surveyed in (Chen2012-dv; Gao2021-xo; Jannach2021-bp)) as a new task for evaluating hybrid models. Early critiquing systems enable the user to provide feedback on recommended items, typically at the feature-level, through (e.g.) suggested pre-designed tweaks (Burke1996-rv) or language-driven dialog systems (Shimazu2002-uq; Thompson2004-jj; Grasch2013-kw). Recent work (Wu2019-qx; Luo2020-uw; Luo2020-qh) explore systems that jointly learn latent representations for users, items, and key-phrases, the latter of which serves as critiques that can influence the recommendations through operations in the latent space. Others explore the interpretation of attributes (Nema2021-ad; Balog2021-cj) to support critiquing. We view hybrid models that associate item sequences with the associated textual context signals as natural extensions to the recent critiquable recommenders. As such, we believe hybrid recommenders offer a promising approach for building critiquing systems, and that it would be instructive to start evaluating such recommenders with critiquing.
3 Method
3.1 Task Definition
We formulate the recommendation task as follows. Assume a dataset consisting of user sequences $S=\{i_{1},i_{2},...,i_{n}\}$ containing items $i\in\mathcal{I}$. Each item $i$ is a tuple $(\mathrm{ID},T)$, where $\mathrm{ID}$ is the unique identifier of the item, and $T=\{t^{1},t^{2},...,t^{m}\}$ is a natural language description of the item consisting of text tokens $t$. Each description $T$ comprises the title, description category, and a short review summary given by the user. The task is to predict $i_{n+1}$ given $S$.
3.2 Flare Model Architecture
(a) Flare Model Architecture.
(b) Flare adapted to incorporate recommendation critiques. Critiques are textual descriptions, incorporated as a separate modality to the Perceiver, to direct the model’s predictions.
(c) Text-ID contrastive alignment loss, aligning textual embeddings with collaborative filtering embeddings. Gray squares are in-batch negative text-item embedding pairs, and yellow squares represent positive pairs.
Figure 1: Sequential recommender model architecture with text augmentation. Text tokens are encoded with a frozen (blue) LLM (mT5), and the associated embeddings are resampled with the Perceiver. Parameters in green boxes are learned using MLM.
Flare augments the Bert4Rec architecture (bert4rec) with item context features (Figure 1). Sequential recommendation models like Bert4Rec use masked language modeling (MLM) with a cloze learning task, similar to BERT’s approach for language tokens. During training, a random subset of items in the input sequence is replaced with a special mask token $\langle mask\rangle$. The model’s objective is to predict these masked items based on their surrounding context. For example, an input sequence of items $[ID_{1},ID_{2},ID_{3},ID_{4},ID_{5}]$ could be transformed through random masking into $[ID_{1},\langle mask\rangle_{1},ID_{3},\langle mask\rangle_{2},ID_{5}]$. The labels for this masked sequence would then be $\langle mask\rangle_{1}=ID_{2}$, $\langle mask\rangle_{2}=ID_{4}$.
Following the item embedding lookup, the inputs to the initial transformer layer can be represented as:
$[e_{1},e_{\langle mask\rangle},e_{3},e_{\langle mask\rangle},e_{5}]$
where $e_{i}$ represents the item embedding for item $ID_{i}$, and $e_{\langle mask\rangle}$ denotes the mask embedding.
The transformer layers process the masked sequence, and the outputs corresponding to the masked positions are fed into a softmax function over the item vocabulary. The model’s loss is calculated as the average negative log-likelihood over the masked positions:
$\mathcal{L}_{MLM}=\frac{1}{S^{m}}\sum_{ID_{m}\in S^{m}}P(ID_{m}=ID_{m}^{*}|S^{\prime})$ (1)
where $S^{m}$ is the set of masked items, $ID_{m}^{*}$ is the actual item corresponding to the masked item $ID_{m}$, $S^{\prime}$ is the input sequence with masking and $P(\cdot)$ represents the output of the softmax layer given the masked input.
To incorporate item contextual information, we augment the item embeddings in the masked sequence as
$[c_{1},e_{\langle mask\rangle},c_{3},e_{\langle mask\rangle},c_{5}]$
where $c_{i}$ is derived from a combination of the item embedding $e_{i}$ and the item context embedding $t_{i}$ representing item metadata. $t_{i}$ is generated using a pre-trained and frozen LLM encoder and a perceiver (perceiver) to encode the LLM output embeddings into a fixed length. The resulting output is concatenated with the item embedding and projected to the modified item embedding $c_{i}$ at the transformer input. This allows the model to learn masked item prediction with text as side information, which we show improves the performance on recommendation tasks. In our experiments, we use the mT5 Base model (mt5) to encode text. The output of the perceiver is then linearly projected to match the item embedding dimensions and then vector summed with the item embeddings.
The use of a perceiver network provides certain key advantages in this context. It encodes a variable length input sequence of embeddings to a fixed length output sequence of embeddings. By concatenating (in the sequence dimension) different modality types at its input, the network can learn a joint encoding. Thus the perceiver offers a natural method of joining multiple modalities of text data, as described later for critiquing.
3.3 ID-Text Contrastive Learning
We implement an auxiliary contrastive loss to align item ID embeddings with item context embeddings as illustrated in Figure 1c. This approach is motivated by the observation that the less frequent, or “tail” items, which have limited representation in training data, can rely instead on their contextual representations. Conversely, frequently appearing items can benefit from collaborative filtering signals that can be learnt from training a sequential recommender. Aligning ID and context embeddings embeds both the collaborative filtering (CF) and contextual signals into the same latent space, allowing context to potentially fill in for gaps in CF. More precisely, we employ the InfoNCE loss function (infoNCE) as detailed below:
$\mathcal{L}_{c}=-\frac{1}{N}\sum^{N}_{i=1}\frac{exp(\frac{e_{i}\cdot t_{i}}{\tau}-m)}{exp(\frac{e_{i}\cdot t_{i}}{\tau}-m)+\sum^{N}_{j=1,j\neq i}exp(\frac{e_{i}\cdot t_{j}}{\tau})}$ (2)
where $\tau$ is the softmax temperature, and we also include a softmax additive margin term $m$ to each positive pair to increase the separation between positive and negative pairs, as proposed by (yang2019improving).
3.4 Critiquing
A core component of hybrid recommenders is their ability to effectively utilize both textual content semantics and collaborative filtering information. However, bridging the gap between these two modalities poses a challenge, potentially leading to a situation where the downstream sequential recommender fails to fully exploit the information available in each.
We propose a novel masking and evaluation paradigm utilizing critiquing to assess whether hybrid recommenders can appropriately leverage both modalities. This consists of adding a critique in the form of a textual product category $S$ to each item $i$ in the user’s sequence $S$, providing a broad hint to direct the model to the next recommended item. That is, each item $i$ in the user sequence $S$ becomes $i=(ID,T,S)$, where $S$ is the critique string. For masked locations during training and inference, we only mask the $ID$ and item description $T$, leaving $S$ unmasked, which the model can utilize as a hint towards the target item (i.e., $i=([\langle mask\rangle,\langle mask\rangle,S)$).
The key to this task is that hints are textual, therefore requiring recommenders to adapt to a text modality by relating the critique to the other signals in the user’s sequence. Note that this task formulation has applications to many real-life uses of recommender systems, such as refining recommendations based on the user’s natural language query text or the webpage subsection the user is currently navigating. In essence, critiquing strings can be used as a vehicle to steer a model predictions to incorporate the user’s current context.
We incorporate critiquing text in Flare as shown in Figure 1. Critique text is a distinct type of input modality, separate from the item contextual text. This is done by separately encoding each modality, with distinct type embeddings, and then using the Perceiver network to merge both text modalities together. Given an item $i_{k}=(ID_{k},T_{k},S_{k})$ consisting of an $ID_{k}$, natural language description $T_{k}$ and critique $S_{k}$, we compute the modified item contextual representation $t_{k}$ and augmented item representation $c_{k}$ with critiquing text as
$T^{\prime}_{k}$ $=LLM(T_{k})+\theta_{T}$ (3)
$S^{\prime}_{k}$ $=LLM(S_{k})+\theta_{S}$ (4)
$t_{k}$ $=Linear(Perceiver([S^{{}^{\prime}}_{k};T^{{}^{\prime}}_{k}]))$ (5)
$c_{k}$ $=t_{k}+e_{k}$ (6)
where $\theta_{T}$ and $\theta_{C}$ are type embeddings to denote text and critiquing as distinct types, “;” denotes concatenation, and “$+$” denotes vector addition, and $e_{k}$ is the embedding for item $ID_{k}$. When masking items for training or during inference, we replace $T^{\prime}_{k}$ and $e_{k}$ at the desired positions with $\langle mask\rangle$ embeddings, allowing the model to utilize $S^{\prime}$ as a hint. By treating critique text as a separate input modality, we hypothesize that the model can learn the connection between critique text and item context. Note that critiquing is applied with Flare through an optional additional input to the Perceiver model, which can be omitted.
3.5 Training
We train our model using MLM loss by masking items in the user’s sequence with 15% probability, ensuring that at least one item is masked. For the total loss, we use a weighted sum of MLM loss and the ID-Text contrastive loss:
$\mathcal{L}_{total}=\alpha\mathcal{L}_{MLM}+(1-\alpha)\mathcal{L}_{c}$ (7)
controlled by the hyper-parameter $\alpha$, which we found setting $\alpha=0.5$ yielded optimal results. For the contrastive loss, we utilize a softmax temperature of $\tau=0.3$ and softmax additive margin term of $m=0.2$.
4 Results
This section investigates the following research questions:
  • RQ1: Do the previously reported Bert4Rec baseline models have potential for further performance gains through tuning?
  • RQ2: Can simple models like Flare scale to large item count datasets?
  • RQ3: Can Flare relate textual critique commands to sequential user signals?
  • RQ4: Can Flare adapt to critique commands that are out of sequence?
4.1 Datasets
Dataset # Users # Items Avg. Length
Games 55 132 17 389 8.99
Office 101 357 27 932 7.87
Scientific 11 011 5 327 6.97
Music 27 525 10 611 8.4
Arts 56 110 22 855 8.77
Pets 47 569 37 970 8.84
Clothing 1 219 592 376 843 9.2
Table 1: Amazon product reviews dataset statistics
We use datasets from Amazon Product Reviews (ni-etal) to evaluate our model. Table 1 shows the statistics [^1] We observed that the statistics from the downloaded Recformer datasets differ from the statistics reported in the Recformer paper, most notably the Games dataset. for the datasets.
Amazon Product Reviews
The Amazon Product Reviews dataset contains 233.1 million reviews from May 1996 to October 2018, covering 29 categories ranging from Office Supplies to Video Games. Each review includes item features (title, description, category, price) and review features (text, timestamp, score, summary). User sequences can be created by sorting each user review by timestamp, creating a sequential recommendation task with the goal of predicting what the user will review next.
We compare our results to Recformer (recformer). To enable fair comparison, we use the six dataset splits provided in Recformer: “Video Games”, “Office Supplies”, “Industrial and Scientific”, “Arts, Crafts and Sewing”, “Pet Supplies” and “Musical Instruments.” Each item has at least five reviews, and every user has given at least five reviews. We use the same preprocessing procedure as Recformer, which includes trimming the user sequences to a max length of 51; training on all sequences, regardless of length; and, evaluating on only sequences with 4 or more items for validation and test. Recformer associates and trains each item with the category, title, and brand features.
In addition to the six “small” datasets above, we also evaluate on one medium-sized dataset, “Clothing, Shoes and Jewelry,” which contains 378k items. The preprocessing procedure is the same as before, but rather than trim, we filter all user sequences with more than 50 items and items with a missing title. We associate and train each item to the category and item features. Notably, this dataset is approximately 9x larger than the largest small dataset. Statistics for this dataset too can be found in Table 1.
4.2 Experimental Setup
Baselines
We compare Flare Text + ID against two classes of state-of-the-art sequential recommenders:
  1. 1.
    ID only models, like Bert4Rec (bert4rec), which utilizes a transformer-based approach with BERT-style (bert) masked language-model training for item sequences. We present updated baseline results after extensive tuning, which we call Flare ID. We emphasize that Flare ID is a Bert4Rec model, with our hyper-parameter tuning.
  2. 2.
    Text-only models, specifically Recformer (recformer), which departs from using explicit item ID representations and proposes a text-based representation using learned key-value pairs.
  3. 3.
    Hybrid Text + ID models, specifically S3Rec (s3rec), which utilizes self-supervised pre-training tasks to align items and attributes through mutual information maximization.
Evaluation Settings
Params Games Office Scient. Music Arts Pets
# Layers 2 2 4 8 2 2
# Heads 2 2 16 8 8 16
Model Dims 64 64 512 256 768 1024
Hidden Dims 256 256 2048 1024 3072 4096
Learning Rate 1e-3 1e-3 1e-4 1e-5 1e-4 1e-4
Batch 1 1 16 16 16 32
Total Steps
50K 50K 50K 50K 50K 10K
Table 2: Hyper-Parameters for Flare ID (tuned Bert4Rec)
Params Games Office Scient. Music Arts Pets
# Layers 8 2 2 2 4 2
# Heads 16 4 8 2 4 8
Model Dims 1024 768 256 1024 512 256
Hidden Dims 4096 3072 1024 4096 2048 1024
Learning Rate 1e-4 1e-4 1e-4 1e-5 1e-4 1e-4
Batch 2 16 8 2 8 16
Total Steps
5K 5K 5 25K 10K 25K
Perceiver #Heads
16 16 8 8 2 8
Perceiver #Layers
8 2 2 8 2 2
Perceiver #Latents
2 8 2 4 2 2
Weight Decay
1e-3 1e-2 1e-3 1e-3 1e-3 1e-3
Table 3: Hyper-Parameters for for Flare T+ID
Following prior work (recformer; lc-rec; bert4rec), we adopt the leave-one-out strategy for evaluation. We reserve each user’s penultimate and last items for validation and test, respectively, using the rest of the user sequence for train.
Since training sequences can be of varying length (Table 1), we pack multiple training sequences into a single input example such that each packed example is as close to the same size as possible. Then, during training, we restrict attention to just the items within the same sequence. This allows us to maximize the throughput of our hardware and avoid wasteful computation due to excessive padding tokens. As a result, the effective batch size during training can vary from step to step, depending on the length of the sequences.
For our experiments, we use the Adam optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.99$. We also performed a hyper-parameter sweep on: the number of the transformer’s layers, heads, model and hidden dimensions, learning rate, batch size, total step count, weight decay, usage of contrastive loss, and number of Perceiver layers, and number of heads and latents for the Perceiver network. Tables 4.2 and 4.2 summarize the best model configurations for each dataset using Recall@10.
4.3 RQ1: Flare Performance on Existing Benchmarks
User Input History
Seq. Title Category
1 Spectrum Diversified Scroll Mail and Letter Organizer, Wall Mount, Black Office Products - Office& School Supplies - Forms, Recordkeeping & Money Handling - Key Cabinets, Racks & Holders
2 Spectrum Diversified Scroll Mail and Letter Organizer, Wall Mount, Black Office Products - Office & School Supplies - Forms, Recordkeeping & Money Handling - Key Cabinets, Racks & Holders
3 Canon PGI-250/CLI-251 Combo Pack Office Products - Office & School Supplies - Printer Ink & Toner - Inkjet Printer Ink
4 Arrow Fastener 506 Genuine T50 3/8-Inch Staples, 1250-Pack Office Products - Office & School Supplies - Staplers & Punches - Staples
Recommendations using Bert4Rec
Rank Title Category
1 Scotch Thermal Laminator 2 Roller System (TL901C) Office Products - Office Electronics - Presentation Products - Laminators
2 Spectrum Diversified Scroll Mail and Letter Organizer, Wall Mount, Black Office Products - Office & School Supplies - Forms, Recordkeeping & Money Handling - Key Cabinets, Racks & Holders
3 Scotch Thermal Laminating Pouches, 8.9 x 11.4-Inches, 3 mil thick, 20-Pack (TP3854-20), Clear Office Products - Office Electronics - Presentation Products - Laminating Supplies
4 Quartet Dry Erase Board, Whiteboard / White Board, Magnetic, 17” x 23”, Assorted Frame Colors - Color Will Vary, 1 Board (MWDW1723M) Office Products - Office & School Supplies - Presentation Boards - Dry Erase Boards
5 Sakura Pigma 30062 Micron Blister Card Ink Pen Set, Black, Ass’t Point Sizes 6CT Set Office Products - Office & School Supplies - Writing & Correction Supplies - Pens & Refills - Calligraphy Pens
Recommendations using Flare (item ID + textual context)
Rank Title Category
1 Sharpie 37001 Permanent Markers, Ultra Fine Point, Black, 12 Count Office Products - Office & School Supplies - Writing & Correction Supplies - Markers & Highlighters - Permanent Markers
2 Canon PGI-250XL Pigment Black Ink, Twin Value Pack, Compatible to MG5520, MG6620, MG5420, … Office Products - Office & School Supplies - Printer Ink & Toner - Inkjet Printer Ink
3 Arrow Fastener 506 Genuine T50 3/8-Inch Staples, 1250-Pack Office Products - Office & School Supplies - Staplers & Punches - Staples
4. AT&T CL2940 Corded Phone with Caller ID/Call waiting, Speakerphone, XL Tilt Display, … Office Products - Office Electronics - Telephones & Accessories - Landline Phones - Corded Telephones
5 Pilot G2 Retractable Premium Gel Ink Roller Ball Pens Fine Pt (.7) Dozen Box Red ; Retractable, Refillable & Premium … Office Products - Office & School Supplies - Writing & Correction Supplies - Pens & Refills - Rollerball Pens - Gel Ink Rollerball Pens
Recommendations using Flare, with label’s category as critique: Office Products - Office & School Supplies - Staplers & Punches
Rank Title Category
1 T50 Heavy Duty Staples [Set of 2] Office Products - Office & School Supplies - Staplers & Punches
2 Onotio Heavy Duty 100 Sheet High Capacity Office Desk Stapler with 1000 Box Staples Office Products - Office & School Supplies - Staplers & Punches
3 Swingline Stapler, Optima 70, Desktop Stapler, 70 Sheet Capacity, Reduced Effort, Half Strip, Silver (87875) Office Products - Office & School Supplies - Staplers & Punches
4 1 X Flat Clinch Staples Mini Box of 1000 by MAX No.10 Office Products - Office & School Supplies - Staplers & Punches
5 Swingline Automatic Stapler, Breeze, 20 Sheet Capacity, Battery Powered, Color Selected For You (S7042131) Office Products - Office & School Supplies - Staplers & Punches
Table 10: Example input sequence and recommendations from Bert4Rec, Flare, and Flare with a category-based critiquing input. The critique is the category of the held-out labeled item.
User Input History
Seq. Title Category
1 Spectrum Diversified Scroll Mail and Letter Organizer, Wall Mount, Black Office Products - Office& School Supplies - Forms, Recordkeeping & Money Handling - Key Cabinets, Racks & Holders
2 Spectrum Diversified Scroll Mail and Letter Organizer, Wall Mount, Black Office Products - Office & School Supplies - Forms, Recordkeeping & Money Handling - Key Cabinets, Racks & Holders
3 Canon PGI-250/CLI-251 Combo Pack Office Products - Office & School Supplies - Printer Ink & Toner - Inkjet Printer Ink
4 Arrow Fastener 506 Genuine T50 3/8-Inch Staples, 1250-Pack Office Products - Office & School Supplies - Staplers & Punches - Staples
Recommendations with alternate in-history category as critique: Forms, Recordkeeping & Money Handling - Key Cabinets, Racks & Holders
Rank Title Category
1 Spectrum Diversified Scroll Mail and Letter Organizer, Wall Mount, Black Office Products - Office & School Supplies - Forms, Recordkeeping & Money Handling - Key Cabinets, Racks & Holders
2 Lucky Line Designer Pattern Key Identifier Caps, 4 Pack (16304) Office Products - Office & School Supplies - Forms, Recordkeeping & Money Handling - Key Cabinets, Racks & Holders
3 KeyGuard SL-8548-U Dual Access Combination Key Cabinet With Chrome 4-Dial Combi-Cam Ultra - 48 Hook Office Products - Office & School Supplies - Forms, Recordkeeping & Money Handling - Key Cabinets, Racks & Holders - Cabinets
4 Master Lock Small Lock Box with 20 Key Capacity, 7131D, Beige Office Products - Office & School Supplies - Forms, Recordkeeping & Money Handling - Key Cabinets, Racks & Holders - Cabinets
5 Security Large Magnetic Hide-A-Key Holder for Over-Sized Keys - Extra-Strong Magnet Office Products - Office & School Supplies - Forms, Recordkeeping & Money Handling - Key Cabinets, Racks & Holders - Cabinets
Recommendations with outside-history category as critique:
Office Products - Office Electronics - PDAs, Handhelds & Accessories - PDA & Handheld Accessories
Rank Title Category
1 Kyocera Kona Black (Virgin Mobile) Office Products - Office Electronics - PDAs, Handhelds & Accessories - PDA & Handheld Accessories
2 iPad Pro 9.7 Case, Apple iPad Pro 9.7 Folio Case, roocase Dual View Pro PU Leather Folio Stand Case for Apple iPad PRO 9.7-inch (2016), Black - NOT Compatible with iPad PRO 10.5 and 12.9, Black Office Products - Office Electronics - PDAs, Handhelds & Accessories - PDA & Handheld Accessories
3 Saunders 00497 RhinoSkin Aluminum Hardcase for Palm Tungsten TX/T5 Office Products - Office Electronics - PDAs, Handhelds & Accessories - PDA & Handheld Accessories
4 ZAGG Folio Case with Backlit Bluetooth Keyboard for Apple iPad Mini Purple-Black (1st Generation Only) Backlit Keyboard and Case Office Products - Office Electronics - PDAs, Handhelds & Accessories - PDA & Handheld Accessories
5 HQRP Replacement Battery Palm Tungsten E2 PDA + Screwdriver + HQRP Universal Screen Protector Office Products - Office Electronics - PDAs, Handhelds & Accessories - PDA & Handheld Accessories
Recommendations with the label’s broadened category as critique: Office Products - Office & School Supplies
Rank Title Category
1 Canon CLI-8M Magenta Ink Tank Office Products - Office & School Supplies
2 Swingline Stapler, Optima 25, Full Size Desktop Stapler, 25 Sheet Capacity, Reduced Effort, Blue/Gray (66404) Office Products - Office & School Supplies
3 Mead #10 Envelopes, Press-It Seal-It, White, 50/Box (75024) Office Products - Office & School Supplies
4 Columbian #10 Envelopes, Security Tinted,4-1/8” x 9-1/2”, White, 500 Per Box, 2-PACK (CO128) Office Products - Office & School Supplies
5 Five Star Filler Paper, Graph Ruled Paper, 100 Sheets/Pack, 11” x 8-1/2”, Reinforced, Loose Leaf, White (17016) Office Products - Office & School Supplies
Recommendations with a the label’s critique, made more specific: Office Products - Office & School Supplies - Staplers & Punches - Staples
Rank Title Category
1 Arrow Fastener 506 Genuine T50 3/8-Inch Staples, 1250-Pack Office Products - Office & School Supplies - Staplers & Punches - Staples
2 Swingline Staples, S.F. 4, Premium, 1/4” Length, 210/Strip, 5000/Box, 1 Box (35450) Office Products - Office & School Supplies - Staplers & Punches - Staples
3 Swingline Staples, Standard, 1/4” Length, 210/Strip, 5000/Box, 5 Pack (35101) Office Products - Office & School Supplies - Staplers & Punches - Staples
4 Swingline Staples, Standard, 1/4” Length, 210/Strip, 5000/Box, 1 Box (35108) Office Products - Office & School Supplies - Staplers & Punches - Staples
5 Arrow Fastener 224 Genuine P22 1/4-Inch Staples, 5,050-Pack Office Products - Office & School Supplies - Staplers & Punches - Staples
Table 11: Example input sequence and recommendations from Flare with various alternative category-based critiques.